(no subject)

John Shields (jshields@ksera)
Tue, 6 Jul 1999 00:00:24 -0500

Shift Summary for eve shift, 07/05/99

On Shift: J. Shields, Bob T.

Hours of beam available at all intensities: 0.75
Hours of beam available at nominal intensity: 0.75
Hours of E832 data (special runs not included): 0.75
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

problems w. interlocks, NM2EU, & tevatron refrigeration.

========================================================================
Runs ended this shift:
Run 13817 - 12.6M events - BEAM_832

Tapes brought to FCC this shift:
(including tapes for online split)

none.

========================================================================
Accesses during shift: 2

20:00 replaced power supply & took laser distance reading

20:47 return downstairs to retrieve item left behind during prev. access

========================================================================
Overview of shift:

Went on access, "blue-iced" alcohol bubbler, watched Sydney struggle to
to trouble-shoot the tape drives,and waited in vain for beam.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13817 16:13 Control Room calls and says they're going to swap
NM2EU power supply. Beam goes away.

18:04 For lack of other diversions, we stop the run.

13818 18:16 Run 13816 begins. All we need now is beam.

18:33 put fresh online-split tapes QKLA06 & QKSM04
into drives tps2d4 & tps3d2, respectively.

19:30 blue ice around alcohol bubbler swapped

19:52 Chan 1980 on CsI trips "val alarm"

20:00 Bob T. and JS go on access to swap MU3V power supply
and to do a laser distance reading

20:45 Return from access. We swapped the MU3V with an unused
FLUKE 1570 HV Supply for the Halo Counters in Station #1.
Front Panel setting of -2270V only achieves -2250V of
actual output, according DVM.

20:46 JS realizes that he oh-so-cleverly left the post-it note
with laser readings downstairs. Oops.

20:47 JS, Bob T. do an encore performance of their access.
The offending post-it note is successfully
liberated. JS blushes.

20:50 return from access. Laser distance readings as follows:
-1.3: 3.9932
-6.0: 38.5451
+2.5: 44.6623

21:07 blue ice around alcohol bubbler swapped. Again.

21:55 OLS problem tps2d4/ktev3, tape QKLA06 (same one that
was requested at 18:30). Online-split gurus paged.

21:55, 10 sec. Sydney walks into Ktev, right on cue. Shift
crew is appropriately impressed.

22:18 blue ice around alcohol bubbler swapped. Again.

22:20 Sydney decides to kill OLS job on QKLA06 & cleans drive

22:50 cleaning the tape drive doesn't help-- tape fails again.
Sydney decides to try a different tape.

23:07 run 13818 aborts! Sydney suspects provider on
ktev3/drive2 drops out, possibly due to tape error?

23:30 We extract ktev3 data tape QKB031 to see if there is, in
fact, any data on it. We load it into a ktev3 OLS drive
to run ktscan, but it will not mount *or* eject!!
Sydney does not look pleased.


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

be sure to keep swapping the blue ice around the alcohol bubbler.